# Verifix: Post-Training Correction to Improve Label Noise Robustness with Verified Samples

### v1

@@ 4.3.4 Projection Matrix :
5 Experiments
5.1 Toy Illustration
(a) Clean Dataset.
(b) Corrupt Dataset
(c) Verifix applied to (b).
Figure 4: Decision boundaries for a network trained the 2D spiral data (a) with full Clean data, (b) with $10\%$ corrupt data and (c) label correction algorithm applied to network trained on corrupt data.
In order to demonstrate our algorithm, we trained a network for binary classification task on clean training data and data where $10\%$ labels were flipped (corrupted) from 2D spiral dataset [42]. Figure 3(a) and 3(b) show the decision boundary of the model for the former and later cases respectively. We observe that, with label corruption, the test accuracy of the model drops by $4.8\%$ from the model trained without any label corruption. The model trained on noisy data(Figure 3(b)) has a more complex and rough decision boundary which indicates memorization of the corrupt data points leading to a loss in generalization. When Verifix is applied to the model trained on corrupt data (Figure 3(c)), we observe the decision boundary smoothness and we recover the generalization performance.
5.2 Synthetic Noise
We evaluate our algorithm on the CIFAR[23] dataset with synthetic noise, using modified ResNet18 [15] and VGG11 (with BatchNorm) [37] architectures. We introduce random label corruption to $x\%$ of the training data $\mathcal{D}_{Tr}$, creating a corrupted dataset $\mathcal{D}^{cpt}_{Tr}$. This corrupted dataset is further split into Validation ($10\%$) and Training ($90\%$) sets. A subset of the corrupt dataset with correctly labeled samples is called a clean subset of the corrupt dataset and is denoted by $\mathcal{D}^{cln}_{Tr}$. We use the Validation set to tune the hyperparameter $\alpha$ and report test accuracy results in Table 1. Training details are provided in Appendix 0.B.
5.2.1 Benchmarks :
We compare the following models in Table 1: (1) Ideal - Model trained with Stochastic Gradient Descent (SGD) on $\mathcal{D}_{Tr}$, entire dataset with clean labels; (2) Retrain - Model trained with SGD on $\mathcal{D}^{cln}_{Tr}$, clean partition of corrupt dataset; (3) Vanilla - Model trained with SGD on $\mathcal{D}^{cpt}_{Tr}$, entire corrupt dataset; (4)SAM - Model trained with Sharpness-Aware Minimization (SAM) [9] on $\mathcal{D}^{cpt}_{Tr}$; (5)Finetune - Vanilla model further finetuned on a small, randomly selected clean subset of $\mathcal{D}^{cln}_{Tr}$; (6) VerifixVanilla - Our method applied to the (pre-trained) Vanilla model; and (7) VerifixSAM - Our method applied to the SAM model.
5.2.2 Results :
Our results in Table 1 show the Retrain models perform closely match the Ideal model, indicating that mildly reducing dataset size has minimal impact on model generalization [13]. In contrast, the Vanilla model experiences a significant drop in test accuracy, especially at higher noise levels. We observe a $15-20\%$ drop with $25\%$ corruption level for both CIFAR10 and CIFAR100. This highlights the severe impact of label noise on performance. Finetuning the Vanilla model on a small set of trusted samples (1000 for 50 epochs) consistently improves performance. Notably, finetuning on just 1000 samples can boost performance by approximately $8\%$ for a VGG11 model trained on CIFAR10 with $25\%$ corruption (see Finetune in Table 1). Applying our method to the Vanilla model (VerifixVanilla) leads to significant improvements on both CIFAR10 and CIFAR100 datasets. With a $25\%$ corruption rate, we observe gains of $11.19\%$ (VGG11) and $6.97\%$ (ResNet18) on CIFAR10, and $7.78\%$ (VGG11) and $3.5\%$ (ResNet18) on CIFAR100.
@@ 5.2.2 Results :
Table 1: Test Accuracy for label noise removal on CIFAR10 and CIFAR100 datasets trained on VGG11 and ResNet18 architectures with $10\%$ and $25\%$ label corruption. We report the mean and standard deviation across three randomly chosen seeds. Verifix and Finetune methods use 1000 trusted samples. See Appendix 0.C for a lower level of synthetic noise.
Method VGG11_BN ResNet18
Percentage Mislabeled Samples Percentage Mislabeled Samples
10% 25% 10% 25%
CIFAR10 Ideal $91.18\pm 0.17$ $94.23\pm 0.06$
Retrain $90.73\pm 0.15$ $89.60\pm 0.21$ $93.77\pm 0.25$ $93.08\pm 0.21$
Vanilla $85.01\pm 0.38$ $72.19\pm 0.33$ $88.66\pm 0.38$ $78.37\pm 0.18$
SAM $\textbf{89.96}\pm\textbf{0.1}$ $87.05\pm 0.28$ $90.80\pm 0.13$ $83.76\pm 0.11$
Finetune $87.03\pm 0.42$ $81.82\pm 0.25$ $89.49\pm 0.2$ $82.82\pm 0.70$
VerfixVanilla $87.78\pm 0.18$ $83.38\pm 0.47$ $90.46\pm 0.10$ $86.96\pm 0.14$
VerfixSAM $89.95\pm 0.11$ $\textbf{87.17}\pm\textbf{0.38}$ $\textbf{91.42}\pm\textbf{0.14}$ $\textbf{87.74}\pm\textbf{0.38}$
CIFAR100 Ideal $67.44\pm 0.18$ $75.34\pm 0.06$
Retrain $65.99\pm 0.43$ $63.73\pm 0.52$ $73.89\pm 0.16$ $72.13\pm 0.36$
Vanilla $60.66\pm 0.23$ $49.03\pm 0.69$ $68.16\pm 0.40$ $57.60\pm 0.17$
SAM $62.58\pm 0.52$ $54.31\pm 0.21$ $69.64\pm 0.18$ $58.82\pm 0.76$
Finetune $60.68\pm 0.23$ $49.01\pm 0.70$ $68.42\pm 0.53$ $58.67\pm 0.17$
VerfixVanilla $63.10\pm 0.40$ $56.81\pm 0.37$ $69.16\pm 0.16$ $61.42\pm 0.46$
VerfixSAM $\textbf{63.80}\pm\textbf{0.50}$ $\textbf{58.33}\pm\textbf{0.11}$ $\textbf{70.21}\pm\textbf{0.16}$ $\textbf{62.19}\pm\textbf{0.66}$
We compare our approach to SAM [9], which is a training technique more robust to label noise but requires double the resources of standard SGD. While SAM s more robust to label noise over the Vanilla model, our method applied to the Vanilla model (VerifixVanilla in Table 3) outperforms SAM in most cases with $25\%$ corruption and remains competitive at $10\%$ corruption. The only exception is VGG11 on CIFAR10, where SAM shows slightly higher gains ($2.18\%$ at $10\%$ corruption, $3.67\%$ at $25\%$). Importantly, VerfixVanilla demonstrates an average improvement of $0.58\%$ over SAM, highlighting its effectiveness while requiring significantly fewer computational resources. Additionally, applying our method to SAM-trained models (VerifixSAM in Table 1) further boosts accuracy up to $4.02\%$, achieving the best results in Table 1. This highlights that our method is complementary to existing noise-robust training techniques, offering additional generalization improvements.
Figure 5: Effect of sampling strategy on CIFAR100 dataset for model trained with $25\%$ corrupt labels.
Figure 6: Effect of trusted sample size on CIFAR100 dataset for model trained with $25\%$ corrupt labels.
Figure 5: Effect of sampling strategy on CIFAR100 dataset for model trained with $25\%$ corrupt labels.
Figure 6: Effect of trusted sample size on CIFAR100 dataset for model trained with $25\%$ corrupt labels.
Figure 7: Effect of $\alpha$ for model trained with $25\%$ corrupt labels.
5.3 Analyses
5.3.1 Selecting samples with low curvature improves generalization :
Our approach uses a few trusted samples from the clean portion of the training dataset ($\mathcal{D}^{cln}_{Tr}$). We explored six different sampling strategies in Figure 7. The Expert strategy provides 1000 correctly labeled samples from $\mathcal{D}^{cln}_{Tr}$, requiring access to ground-truth labels. Alternatively, we use curvature scores (Equation 5) to identify likely clean samples from the corrupted dataset $\mathcal{D}^{cpt}_{Tr}$. We sort samples by ascending curvature and select 1000 based on percentile scores (0%, 10%, 25%, 50% strategy) or random sampling from the lower 50% percentile samples (Uniform strategy). By xsampling strategy, we mean selecting 1000 images after the $x^{th}$ percentile of low-curvature samples. Interestingly, selecting the 1000 lowest curvature samples (0% sampling) consistently outperforms other strategies, suggesting these samples are most beneficial for our method. For CIFAR100 dataset with 0% sampling, we achieve $56.85\%$ and $61.1\%$ test accuracy on VGG11 and ResNet18 model, respectively. We observe that performance decreases as we sample from higher curvature percentiles, with steeper declines for VGG11. Our experiments show an curvature based sampling improves the generalization up to $1.06\%$ for 0% strategy when compared to Expert. Additionally, the Expert sampling strategy gives similar performance to Unifrom strategy ($\sim 55.8\%$ for VGG11 and $\sim 60.55\%$ for ResNet18 on CIFAR100).
5.3.2 Increasing Trusted Sample Size yields diminishing returns in performance :
We study how the number of trusted samples impacts our method’s performance and computational cost. Since representation calculations and SVD are performed on the trusted set, the sample size directly affects compute requirements. Figure 7 shows how test accuracy varies with the number of trusted samples. We observe that increasing the sample size improves accuracy up to 1000 samples, after which there are no significant gains. Therefore, we use 1000 samples in our experiments to balance performance and computational efficiency.
5.3.3 Finding the optimal scaling hyperparameter ($\alpha$) is key to maximizing performance gains :
Figure 7 demonstrates how the scaling factor ($\alpha$) influences test accuracy for models trained on datasets with $25\%$ label corruption. We observe that increasing $\alpha$ generally improves test accuracy up to a certain point, after which accuracy gradually declines. Importantly, we find that $\alpha=30000$ consistently yields strong results across different datasets, models, and corruption levels.
Figure 8: Compute comparisons for several methods for ResNet18 model trained on CIFAR100 dataset.
[…]
Figure 8: Compute comparisons for several methods for ResNet18 model trained on CIFAR100 dataset.
5.3.4 Verifix achieves SAM-level performance at a fraction of the computational cost :
We analytically evaluate the computational cost of different methods in Figure 8. In our evaluations, we only consider the convolution and linear layers and ignore the compute for normalization, pooling, and non-linearities in the network. Compute of forward pass for linear layer is $n_{sample}f_{in}f_{out}$, where $n_{sample}$ is the number of samples, $f_{in}$ is the input dimension of linear layer and $f_{out}$ is the output dimension. For a convolutional layer with a kernel of size $C_{in}\times k\times k\times C_{out}$ producing an output map of size $h_{o}\times w_{o}$ requires compute of $n_{sample}h_{o}w_{o}C_{in}k^{2}C_{out}$ for the forward pass. The backward pass cost is approximated as twice the forward pass cost. Note, here $n_{samples}$ is the effective number of samples seen during the training or finetuning. If a model is trained for 50 epochs with 1000 samples in each epoch, then the value for $n_{sample}$ is 50000. This explains the analytical compute costs for Finetune, Retrain, and Ideal baselines, which use standard SGD. SAM, requiring two forward and backward passes per iteration, has its compute cost approximated by doubling the SGD cost.
We decompose Verifix’s computation into five steps: Trusted sample Estimation, Representation Sampling, SVD, Space Estimation, and Weight Projection. Trusted Sample Estimation involves a few gradient computation epochs ( $\sim 10$) over $\mathcal{D}_{Tr}^{cpt}$ using curvature estimation [10]. Representation Sampling requires a forward pass on the trusted sample set. SVD cost for an m×n matrix (where $m>n$) is approximately $mn^{2}$ [6]. Space Estimation and weight projection primarily involve matrix multiplications, where the cost for multiplying an $m\times n$ matrix with an $n\times p$ matrix is $mnp$. This analytical framework, including the previously discussed forward pass and matrix multiplication costs, explains the compute estimates in Figure 8. The bar plot in Figure 8 displays these analytical compute estimates on a log scale, with bar labels indicating normalized compute values relative to Verifix. Verifix demonstrates performance competitive to SAM (see Table 1) while requiring significantly lower computational resources (Figure 8). Particularly, the total compute requirement for Verifix applied to a Vanilla model is less than $\sim 51.5\%$ (calculated as $100\times(34.7+1)/69.4$) of the compute needed by SAM. This highlights Verifix’s efficiency advantage without compromising accuracy.
5.4 Real World Noise
@@
To demonstrate our method’s effectiveness in real-world noise scenarios, we present experiments on Mini-WebVision [17], WebVision1.0 [25], and Clothing1M [46] datasets. Mini-WebVision and WebVision1.0 utilize the InceptionResNetV2 [40] architecture, while Clothing1M employs ResNet50 [15]. Comprehensive training and hyperparameter details are provided in Appendix 0.B. In Table 2, we establish several baselines for these datasets including Vanilla, MixUp [47], MentorMix [17], and SAM [9]. Following previous works, we report validation accuracy for Mini-WebVision and WebVision, and test accuracy for Clothing1M.
5.4.1 Results :
As shown in Table 2, applying Verifix in conjunction with each of the noise-robust training methods consistently yields generalization improvements. These results demonstrate an average improvement of $1.13\%$, highlighting Verifix’s ability to complement existing approaches while maintaining computational efficiency. Importantly, our success with such larger datasets (containing millions of images), underscores the scalability of our method.
6 Discussion and Conclusion
This paper introduces Post-Training Correction, a novel paradigm for addressing label corruption. Unlike data cleaning or noise-robust training methods, our approach directly modifies model weights after training, offering a superior trade-off between generalization and compute. Data cleaning incurs compute overheads to pinpoint mislabeled samples, while noise-robust training can be computationally expensive and may necessitate retraining. Importantly, these existing approaches don’t fully close the generalization gap caused by noisy labels. We propose Verifix, a Post-Training Correction algorithm that complements existing methods and effectively bridges this gap, demonstrating a $1.13\%$ generalization improvement on real-world corrupted datasets while maintaining computational efficiency.
Verifix’s efficiency stems from its focus on a small set of trusted samples, reducing the computation for representation sampling and SVD. The layer-wise independence of SVD enables parallelization for further speed gains. Crucially, Verifix utilizes only correctly labeled samples, which are significantly easier to obtain than identifying corrupt data.
While Verifix offers significant advancements, we believe the Post-Training Correction domain holds the potential for even greater generalization improvements. Our experiments highlight the impact of sampling strategies, suggesting value in exploring computationally efficient methods for selecting trusted data. We invite the research community to address these exciting challenges: developing Post-Training Correction algorithms to close the performance gap, exploring efficient trusted data sampling strategies, and evaluating Verifix on several datasets and network architectures.
Acknowledgement

### v2

@@ 1. $\mathcal{D}_{Trust}$ = get_trusted($\theta_{*}$, $\mathcal{D}_{Tr}$) // Eqn 4
Experiments
Experimental Setup
We evaluate our algorithm on the CIFAR10 and CIFAR100 datasets (Krizhevsky, Hinton et al. 2009) using VGG11 (Simonyan and Zisserman 2015) and ResNet18 (He et al. 2016) models. To simulate various noise scenarios, we introduce synthetic noise to the original training data according to Equations 910, or  11 and subsequently partition the data into a $95\%$ training set ($\mathcal{D}_{Tr}$) and a $5\%$ validation set ($\mathcal{D}_{Val}$). Importantly, we add noise before the train-validation split, mimicking a real-world scenario where noise is present from the outset. Unless otherwise specified, we tune hyperparameters on validation set ($\mathcal{D}_{Val}$) and use the accuracy on the test set ($\mathcal{D}_{Te}$) as the metric to assess generalization performance of the algorithm. In all our tables we report the mean and standard deviation across three randomly chosen seeds in all our experiments. Specific training configurations and hyperparameters are provided in the supplementary material for all the experiments. We explore different type of noise in our evaluations presented in the next subsection.
  Method VGG11_BN ResNet18 Average
Retain Forget Symmetric Noise ( Eqn. 9 ) Symmetric Noise ( Eqn. 9 ) Asymmetric Noise ( Eqn. 10 ) Hierarchical Noise ( Eqn. 11 )
samples samples $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$
CIFAR10 Retrain - - $90.18\pm 0.14$ $89.48\pm 0.04$ $93.21\pm 0.22$ $92.45\pm 0.06$ $93.38\pm 0.07$ $92.75\pm 0.23$ $93.47\pm 0.21$ $93.14\pm 0.23$ $92.26$
Vanilla 00 00 $86.04\pm 0.17$ $76.68\pm 0.48$ $88.55\pm 0.16$ $79.47\pm 0.46$ $88.53\pm 0.34$ $79.79\pm 1.53$ $91.42\pm 0.33$ $86.42\pm 0.38$ $84.61$
Finetune $5000$ 00 $85.47\pm 0.13$ $80.94\pm 0.76$ $88.28\pm 0.30$ $85.16\pm 0.12$ $87.31\pm 0.81$ $82.82\pm 0.98$ $91.42\pm 0.33$ $\textbf{88.23}\pm\textbf{0.73}$ $86.20$
SSD $5000$ $1000$ $86.00\pm 0.21$ $76.77\pm 0.58$ $88.54\pm 0.17$ $79.48\pm 0.50$ $88.52\pm 0.35$ $80.36\pm 1.45$ $91.42\pm 0.33$ $86.39\pm 0.42$ $84.68$
SCRUB $1000$ $200$ $85.88\pm 0.35$ $78.90\pm 0.25$ $89.50\pm 0.17$ $83.77\pm 0.44$ $89.50\pm 0.22$ $83.60\pm 0.14$ $91.65\pm 0.12$ $88.00\pm 0.58$ $86.35$
SAP 00 00 $\textbf{87.25}\pm\textbf{0.16}$ $\textbf{82.27}\pm\textbf{0.15}$ $\textbf{90.12}\pm\textbf{0.11}$ $\textbf{85.49}\pm\textbf{0.39}$ $\textbf{90.03}\pm\textbf{0.25}$ $\textbf{86.32}\pm\textbf{0.66}$ $\textbf{91.87}\pm\textbf{0.22}$ $87.92\pm 0.69$ 87.66
CIFAR100 Retrain - - $65.76\pm 0.23$ $63.66\pm 0.47$ $72.43\pm 0.42$ $71\pm 0.15$ $73.76\pm 0.46$ $71.62\pm 0.47$ $72.73\pm 0.19$ $71.16\pm 0.39$ $70.26$
Vanilla 00 00 $60.41\pm 0.14$ $50.64\pm 0.60$ $65.84\pm 0.33$ $54.75\pm 0.45$ $72.98\pm 0.13$ $61.86\pm 0.60$ $67.45\pm 0.12$ $57.82\pm 0.21$ $61.47$
Finetune $5000$ 00 $60.26\pm 0.11$ $52.50\pm 0.31$ $65.97\pm 0.35$ $57.33\pm 0.40$ $72.98\pm 0.13$ $61.29\pm 1.13$ $67.55\pm 0.15$ $59.98\pm 1.11$ $62.23$
SSD $5000$ $1000$ $60.38\pm 0.16$ $50.62\pm 0.60$ $65.84\pm 0.33$ $54.77\pm 0.42$ $72.99\pm 0.11$ $61.67\pm 0.55$ $67.43\pm 0.21$ $57.83\pm 0.21$ $61.44$
SCRUB $1000$ $200$ $60.93\pm 0.09$ $52.11\pm 0.63$ $\textbf{67.02}\pm\textbf{0.29}$ $57.36\pm 0.40$ $\textbf{73.12}\pm\textbf{0.18}$ $63.37\pm 0.69$ $68.20\pm 0.13$ $60.24\pm 0.33$ $62.79$
SAP 00 00 $\textbf{61.10}\pm\textbf{0.23}$ $\textbf{53.31}\pm\textbf{0.78}$ $66.82\pm 0.17$ $\textbf{58.74}\pm\textbf{0.61}$ $72.92\pm 0.30$ $\textbf{63.57}\pm\textbf{0.49}$ $\textbf{68.24}\pm\textbf{0.17}$ $\textbf{60.76}\pm\textbf{0.50}$ 63.18
Table 1: Test Accuracy for symmetric noise (Equation 9) removal from CIFAR10 and CIFAR100 datasets trained on VGG11 and symmetric noise, asymmetric noise (Equation 10) and hierarchical noise (Equation 11) removal ResNet18 architectures with noise strength $\eta=0.1$ and $\eta=0.25$. Note, the confusing class groups (or set $\mathcal{C}$) for hierarchical noise are cat-dog and truck-automobile for CIFAR10. For CIFAR100 we use Superclasses as the confusing groups. - We select 1000 samples with low loss value from $\mathcal{D}_{Tr}$ as retain samples for our method.
Types of Synthetic Noise
We evaluate our approach under various noise settings, characterized by the noise transition matrix, denoted by $\mathcal{T}$. The noise transition matrix $\mathcal{T}$ is a square matrix of size $K\times K$, where $K$ is the number of classes, which captures the conditional probability distribution of label corruption. The element at the $i^{th}$ row and $j^{th}$ column, denoted by $t_{ij}$, describes the probability of a clean label, $c_{i}$, being flipped to a noisy label $c_{j}$. The degree of noise added is controlled by the parameter $\eta$. We study three different noise setting as described below:
Symmetric Noise:
In symmetric noise (Ghosh, Kumar, and Sastry 2017), labels are flipped to any other class with equal probability. This creates a uniform distribution of errors across all classes. Equation 9 defines the element of the noise transition matrix $\mathcal{T}$ under a symmetric noise setting.
$\begin{split}t_{ij}=\begin{cases}\frac{\eta}{(K-1)}\qquad&i\neq j\\ 1-\eta&i=j\end{cases}\end{split}$ (9)
Asymmetric Noise:
In asymmetric noise (Ghosh, Kumar, and Sastry 2017), the probability of a label being flipped to a specific class varies. The noise transition matrix $\mathcal{T}$ in our experiments is defined by Equation 10. The values for off-diagonal elements are sampled from a uniform distribution $\mathcal{U}$ over the range 0 to $\frac{2\eta}{(K-1)}$. The diagonal elements, $t_{ii}$, are computed to ensure that each row of the matrix sums to 1.
$\vspace{-3mm}\begin{split}t_{ij}=\begin{cases}\epsilon\sim\mathcal{U}(0,\frac{2\eta}{(K-1)})\qquad&i\neq j\\ \\ 1-\sum_{k=1;k\neq i}^{k=K}t_{ik}&i=j\end{cases}\end{split}$ (10)
Hierarchical Noise:
To simulate a practical label noise scenario, we introduce label errors based on class hierarchy (Mukherjee, Garg, and Roy 2024). This approach reflects the intuition that human labelers are more likely to confuse hierarchically closer classes. We define a set of confusing class groups, $\mathcal{C}$, where $g_{i}\in\mathcal{C}$, indicates all the classes which are hierarchically closer to class $i$. Equation 11 represent the noise transition matrix for this case.
$\begin{split}t_{ij}=\begin{cases}\frac{\eta}{|g_{i}|}\quad&i\neq j;\hskip 5.69054ptj\in g_{i}\\ 0&i\neq j;\hskip 5.69054ptj\notin g_{i}\\ 1-\sum_{k=1;k\neq i}^{k=K}t_{ik}&i=j\end{cases}\\ \text{ where }g_{i}\in\mathcal{C};\text{ and }|g_{i}|\text{ number of classes in }g_{i}\end{split}$ (11)
Corrective Machine Unlearning Benchmark
We compare our proposed method, SAP, to several unlearning algorithms under different synthetic label noise listed above. As a reference point, we include a Retrain model trained with Stochastic Gradient Descent (SGD) on the clean data partition, $\mathcal{D}^{cln}_{Tr}$. All unlearning methods are applied to the Vanilla model, which is trained with SGD on the entire corrupt dataset, $\mathcal{D}_{Tr}$. We also include a Finetune baseline, which updates all layers of the Vanilla model using SGD on a small retained dataset, leveraging the idea of catastrophic forgetting as explored in CF-k (Goel et al. 2022). For a strong baseline, we implement SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023), a state-of-the-art unlearning algorithm, and SSD (Foster, Schoepf, and Brintrup 2024), which, like SAP, uses a single update for unlearning.
Many unlearning algorithms require access to both retain (correctly labeled) and forget (misclassified) samples (Goel et al. 2024). However, accurately distinguishing between hard and misclassified samples is challenging (Garg, Ravikumar, and Roy 2024). To address this, we provide all unlearning algorithms with small subsets of retain and forget samples for evaluation. Unlike other methods, SAP does not explicitly use forget samples for model correction and uses the sample loss to identify potential samples for the retain set, bypassing the need for curated subsets. We observed other unlearning algorithms suffered significant performance degradation when using samples with low loss as the retained set.
Results:
Our results in Table 1 show that Retrain models closely match expected generalization performance, indicating that a slight reduction in dataset size has minimal impact on model generalization (Guo, Zhao, and Bai 2022). In contrast, the Vanilla model experiences a significant drop in test accuracy, especially at higher noise levels, with a $15-20\%$ decline observed at $25\%$ corruption level for both CIFAR10 and CIFAR100. This highlights the detrimental impact of label noise on performance. Finetuning the Vanilla model for 50 epochs on a small subset of $\mathcal{D}_{Tr}^{cln}$ (the clean partition of $\mathcal{D}_{Tr}$) containing 5000 randomly selected samples consistently improves performance, with a notable $6\%$ gain for a ResNet18 model on CIFAR10 with $25\%$ symmetric label corruption. Interestingly, SSD-based unlearning fails to provide a consistent generalization improvement, which might be attributed to the small sample size, as observed in previous studies (Goel et al. 2024; Kodge, Saha, and Roy 2024). SAP outperforms all other unlearning methods on both CIFAR-10 and CIFAR-100 datasets, achieving an average improvement of $1.36\%$ and $0.39\%$ compared to the second-best method, respectively. In our experiments, we observed that increasing the number of retain and forget samples to 5000 and 1000, respectively, improved the performance of the SCRUB algorithm, as presented in Table 4 in the supplementary material. However, SCRUB requires extensive hyperparameter tuning (retain batch size, forget batch size, $\alpha$, $\gamma$, number of minimization steps, and number of maximization steps) to counteract the instability of gradient ascent, leading to approximately 675 hyperparameter search iterations compared to our method’s 16. Increasing the number of samples further exacerbates the computational burden for hyperparameter search. In the next subsection, we demonstrate how SAP can enhance the generalization performance of label noise robust learning algorithms.
Noise-Robust Learning Benchmarks
  Method Baseline SAP Improvements
CIFAR10 Vanilla $79.47\pm 0.46$ $85.46\pm 0.41$ 5.99
Logit Clip $82.91\pm 0.32$ $85.99\pm 0.67$ 3.08
MixUp $83.12\pm 0.44$ $86.45\pm 0.52$ 3.33
SAM $83.29\pm 0.28$ $87.29\pm 0.08$ 4.0
MentorMix $89.64\pm 0.32$ $90.51\pm 0.17$ 0.87
                                            Average $83.69$ $87.14$ 3.45
  CIFAR100 Vanilla $54.75\pm 0.45$ $58.69\pm 0.68$ 3.94
Logit Clip $56.41\pm 0.43$ $59.90\pm 0.84$ 3.49
SAM $56.49\pm 0.93$ $59.34\pm 0.76$ 2.85
MixUp $58.25\pm 0.65$ $62.32\pm 0.91$ 4.07
MentorMix $68.53\pm 0.35$ $68.98\pm 0.45$ 0.45
                                            Average $58.89$ $61.85$ 2.98
@@ Results:
Table 2: Generalization benefits of applying our algorithm in synergy with noise-robust training approaches for symmetric noise (Equation 9) removal from CIFAR10 and CIFAR100 datasets trained on ResNet18 architectures with noise strength $\eta=0.25$. We report the mean and standard deviation across three randomly chosen seeds.
We compare our method to several label noise robust learning algorithms: Logit Clip (Wei et al. 2023), Sharpness Aware Minimization (SAM) (Foret et al. 2021), MixUp (Zhang et al. 2017), and MentorMix (Jiang et al. 2020), known for their inherent tolerance to mislabeled data. Table 2 presents the performance of various algorithms on CIFAR-10 and CIFAR-100 datasets using the ResNet18 model under symmetric label noise with $\eta=0.25$. Notably, SAP improves generalization performance by $3.45\%$ and $2.98\%$ on average for CIFAR-10 and CIFAR-100, respectively.
Real World Noisy Benchmark
@@
To demonstrate SAP’s effectiveness in real-world noise scenarios, we present experiments on Mini-WebVision (Jiang et al. 2020) and Clothing1M (Xiao et al. 2015) datasets. Mini-WebVision is trained on the InceptionResNetV2 (IRV2 in Table 3) (Szegedy et al. 2017) architecture, while Clothing1M is trained on ResNet50 (He et al. 2016) and ViT_B_16. Comprehensive training and hyperparameter details are provided in supplementary material.
As shown in Table 3, applying SAP consistently yields generalization improvements for a noisy dataset. These results demonstrate an average improvement of $1.8\%$, highlighting SAP’s ability to scale to large datasets (Clothing1M) and SoTA transformer based models (ViT_B_16).
  Dataset Architecture Vanilla SAP Improvement
Mini-WebVision IRV2 $63.81\pm 0.38$ $64.73\pm 0.53$ 0.92
Clothing1M ResNet50 $67.48\pm 0.64$ $69.64\pm 0.57$ 2.16
Clothing1M ViT_B_16 $69.12\pm 0.45$ $71.43\pm 0.60$ 2.31
                      Average $66.80$ $68.60$ 1.80
Table 3: Generalization benefits of applying our algorithm on real-world noisy dataset like Mini-WebVision (Jiang et al. 2020), and Clothing1M (Xiao et al. 2015) for Vanilla trained InceptionResNetv2, called IRV2 in Table (Szegedy et al. 2017) and ViT_B_16 (Dosovitskiy et al. 2021).
Analyses
Increasing Trusted Sample Size yields diminishing returns:
We study how the number of trusted samples impacts SAP’s performance. Figure 6 shows how test accuracy varies with the number of trusted samples. We observe that increasing the sample size improves accuracy up to 1000 samples, after which there are no significant gains. Since representation calculations and SVD are performed on the trusted set, the sample size directly affects compute requirements. Therefore, we use 1000 samples in our experiments to balance performance and computational efficiency.
Optimal $\alpha$ is key to maximizing performance gains :
Figure 6 demonstrates how the scaling factor ($\alpha$) influences test accuracy for models trained on datasets with $25\%$ label corruption. We observe that increasing $\alpha$ generally improves test accuracy up to a certain point, after which accuracy gradually declines. Importantly, we find that $\alpha=30000$ consistently yields strong results for synthetic noise across different datasets, models, and corruption levels.
Figure 5: Effect of trusted sample size for CIFAR100 dataset with $\eta=25\%$.
Figure 6: Effect of $\alpha$ for model trained on CIFAR100 dataset with $\eta=25\%$.
Discussion and Conclusion
This paper introduces SAP, a corrective machine unlearning algorithm that is competitive with the SoTA SCRUB without requiring explicit access to clean and mislabeled data subsets. This eliminates the need for laborious human curation of training data. SAP performs a single model update using Singular Value Decomposition (SVD) on a small sample set, ensuring computational efficiency. Furthermore, with only one hyperparameter, $\alpha$, and reusable SVD computations across different $\alpha$ values, SAP offers simplicity and efficiency, unlike gradient-ascent based algorithms which often require extensive hyperparameter tuning, sometimes exceeding the computational cost of retraining the model. In summary, SAP is a sample and compute-efficient unlearning technique to address the challenges posed by noisy data.
Acknowledgements